Given the following information, formulate a continuous-review inventory management system. The item is demanded 50 weeks a year. 2. Item cost Order cost Annual holding cost (%) Annual demand Average demand $20 $250 10% df item cost 25,000 500 per week Standard deviation of weekly demand Lead time Service probability 25 per week 1 week 95% Compute the order quantity and reorder point. For the order quantity, please provide at least one step of calculation and the correct answer for full credit. For the reorder point, please provide the formula, at least one step of calculation, and the correct answer for full credit. (10 points)
Answer #1

Annual Demand (D) = 25000

Order cost (S) = 250

Holding cost (H) = 10%*Item cost = 10%*20 = 2

Average weekly demand (d) = 500

Standard deviation (SD) = 25 per week

Lead time (LT) = 1 week

Service probability is 95%. Z value for 95% is 1.645

Order quantity = sqrt(2*D*S/H) = sqrt(2*25000*250/2) = 2500

Reorder point = d*LT + z*SD*sqrt(LT) = 500*1 + 1.645*25*sqrt(1) = 541.125


answered by: ANURANJAN SARSAM